OpenStudy (anonymous):

Two boys in a boat with a small motor are able to travel 2 mph faster with the current than against the current. If they travel with the current from a dock to their campground in 1.5 hours and make the return trip against the current in 2.5 hours, how fast are the boys able to travel in still water?

OpenStudy (anonymous):

let s be the speed simply without the current \[s+2=\frac{d}{1.5}\] \[s-2=\frac{d}{2.5}\] \[-2+\frac{d}{1.5}=2+\frac{d}{2.5}\] \[\frac{10d}{15}=4+\frac{6d}{15}\] \[\frac{4d}{15}=4\] d=15 \[s+2=\frac{15}{1.5}\] \[s=8\] \[s-2=\frac{15}{2.5}\] \[s=8\] so wthout current \[8=\frac{15}{t}\] \[t=\frac{15}{8}\]

OpenStudy (anonymous):

lol what?

OpenStudy (anonymous):

OutKast can you solve another problem for me